# Equilibrium under Time-Inconsistency: A New Existence Theory by Vanishing Entropy Regularization

### 11 Mar 2026 v1

@@ 3 Time-inconsistent Control under Entropy Regularization
$\|\bar{V}^{n}\|_{{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(D_{N_{0}})}\leq C\left(\|\bar{V}^{n}\|_{{\mathcal{C}}^{0,0}(D_{N_{0}+1})}+\|f^{n}\|_{{\mathcal{C}}_{\alpha/2,\alpha}(D_{N_{0}+1})}\right),$ (3.47)
@@
We are now ready to prove Theorem 3.1.
Proof.
Fix a $0<\lambda<\lambda_{0}$ and an arbitary $0<\beta<\alpha$. Lemmas 3.3 and 3.4 show that the mapping $\Phi_{\lambda}$ defined in (3.29) is a continuous mapping from the compact convex set ${\mathcal{M}}_{\lambda}$, equipped with norm $\|\cdot\|_{{\mathcal{C}}^{1,2,\operatorname{\textbf{wg}}}_{\beta/2,\beta}(\mathbb{T}\times\mathbb{R}^{d})}$, to itself. Hence, by Schauder fixed-point theorem (see, e.g., [22, Theorem 5.28]), $\Phi_{\lambda}$ has a fixed point $w\in{\mathcal{M}}_{\lambda}$ such that $w=V^{\pi^{*}}_{\lambda}$ with $\pi^{*}(x,a)=\Gamma_{\lambda}(x,D_{x}w(0,x),a)$. As a result, (3.19) in Lemma 3.3 yields that $V^{\pi^{*}}_{\lambda}$ satisfies (3.6). By taking $t=0$ in (3.19), together with (3.12), we see that $V^{\pi^{*}}_{\lambda}$ also satisfies (3.5). Thanks to Proposition 3.1, $\pi^{*}$ is a regularized equilibrium with the entropy parameter $\lambda$. Moreover, the estimates of $V^{\pi^{*}}_{\lambda}$ in (3.11) are direct consequences from $V^{\pi^{*}}_{\lambda}\in{\mathcal{M}}_{\lambda}$ and Lemma 3.3. ∎
@@ 4 Existence of Equilibrium by Vanishing Entropy Regularization
Finally, in this section, we will turn back to the original time-inconsistent control problem, and our goal is to address the existence of equilibrium by establishing the convergence relationship from the entropy regularized problem to the original one when the entropy parameter $\lambda{\to}0$.
Let us take a sequence $(\lambda_{n})_{n\in\mathbb{N}}$ with $\lambda_{n}\to 0+$. For each $n\in\mathbb{N}$, Theorem 3.1 states the existence of a regularized equilibrium $\pi^{n}$ with entropy parameter $\lambda_{n}$ in Gibbs-form such that $\|V^{\pi^{n}}_{\lambda_{n}}\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}\leq A^{*}$ for all $n\in\mathbb{N}$. Let us further denote $v^{n}(t,x):=V^{\pi^{n}}_{\lambda_{n}}(t,x)$ for each $n$. It then follows from Proposition 3.1 that $(v^{n},\pi^{n})$ satisfies
@@
There exist a subsequence $(v^{n_{k}},\pi^{n_{k}})_{k\in\mathbb{N}}$ and a pair $(v^{\infty},\pi^{\infty})$ satisfying the follows:
@@
    $\lim_{k\to\infty}\|v^{n_{k}}-v^{\infty}\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(D_{N})}=0,$ (4.3)
@@
    $\pi^{\infty}:\mathbb{R}^{d}\mapsto{\mathcal{P}}(U)$ is Borel measurable and $\pi^{n_{k}}$ converges to $\pi^{\infty}$ in the sense that
@@
    for any test function $\phi(x,a):\mathbb{R}^{d}\times U\mapsto\mathbb{R}^{d}$ that is continuous on the control variable $a$ for each $x$ and
    $\int_{\mathbb{R}^{d}}\max_{a}|\phi(x,a)|dx<\infty.$
@@
    $V^{\pi^{\infty}}=v^{\infty}$ and it holds that
    $\sup_{x\in\mathbb{R}^{d}}\|v^{n_{k}}(\cdot,x)\|_{{\mathcal{C}}^{1}_{\alpha/2}(\mathbb{T})},\sup_{x\in\mathbb{R}^{d}}\|v^{\infty}(\cdot,x)\|_{{\mathcal{C}}^{1}_{\alpha/2}(\mathbb{T})}\leq A^{*}.$ (4.5)
@@
where $A^{*}$ is the finite constant stated in Theorem 3.1, and $A_{N}=C(N^{d+1})^{1/p}$ with $C$ being a finite constant independent of $d,p$ and $N$. Let us define for each $N$ that
@@
Due to $\|v^{N-1,n_{k}}-u^{N}\|_{W^{1,2}_{p}(D_{N})}\leq A_{N}A^{*}$, Banach–Alaoglu theorem can be called to yield a subsequence of $(v^{N-1,n_{k}})_{k\in\mathbb{N}}$, denote by $v^{N,n}_{n\in\mathbb{N}}$, such that
@@
As a result, (4.6) and the following hold for the subsequence $(v^{N,n})_{n\in\mathbb{N}}$ and $u^{N}$
@@
By taking the diagonal subsequence $(v^{N,N})_{N\in\mathbb{N}}$, we have that, for each $K\in\mathbb{N}$ and each $l\in\mathbb{N}$, the multi-index $a$ with $0\leq 2l+|a|_{l^{1}}\leq 2$,
$\lim_{N\to\infty}\|v^{N,N}-v^{\infty}\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(D_{K})}=0,$
@@
Moreover, for a generic function $g\in L^{p}(D_{1}(t_{0},x_{0}))$, note that
$\|g\|_{L^{p}(D_{1}(t_{0},x_{0}))}=\sup_{\|\phi\|_{L^{q}(D_{1}(t_{0},x_{0}))}\leq 1}\int_{D_{1}(t_{0},x_{0})}\phi(t,x)g(t,x)dtdx.$
Then, for any test function $\phi$ with $\|\phi\|_{L^{q}(D_{1}(t_{0},x_{0}))}\leq 1$, it follows that
@@
$\leq$ $\liminf_{N\to\infty}\|\partial_{t}D^{a}_{x}v^{N,N}\|_{L^{p}(D_{1}(t_{0},x_{0}))},$
which implies that
$\|\partial_{t}D^{a}_{x}v^{\infty}\|_{L^{p}(D_{1}(t_{0},x_{0}))}\leq\liminf_{N\to\infty}\|\partial_{t}D^{a}_{x}v^{N,N}\|_{L^{p}(D_{1}(t_{0},x_{0}))}\leq A^{*},\quad\forall 0\leq 2l+|a|_{l^{1}}\leq 2.$
That is, $\|v^{\infty}\|_{W^{1,2,\operatorname{\textbf{ul}}}_{p}(\mathbb{T}\times)}\leq A^{*}$.
Part (ii). Next, we consider the associated sequence $(\pi^{N,N})_{N\in\mathbb{N}}$ with respect to the sequence $(v^{N,N})_{N\in\mathbb{N}}$, by Young measure theory (see, e.g., [29, Chapter IV]). A subsequence $(\pi^{n_{k}})_{k\in\mathbb{N}}$ of $(\pi^{N,N})_{N\in\mathbb{N}}$ and a Borel function $\pi^{\infty}:\mathbb{R}^{d}\mapsto{\mathcal{P}}(U)$ exist such that part (i) holds. As a result of the continuity of $b,r$ on $a$, it follows from (4.4) that
[…]
In conclusion, the corresponding subsequence $(v^{n_{k}},\pi^{n_{k}})_{k\in\mathbb{N}}$ and the pair $(v^{\infty},\pi^{\infty})$ indeed satisfy (4.7) and Part (i).
Part (iii). Step 1. To prove $V^{\pi^{\infty}}(t,x)=v^{\infty}(t,x)$, we first show that $v^{n_{k}}$ and $v^{\infty}$ vanish as $t\to\infty$. Indeed, by replacing $w$ with $D_{x}v^{n_{k}}(0,x)$ and $V^{\pi}_{\lambda}$ with $v^{n_{k}}$ in (3.20), (3.25) and (3.26), we can see that, for any $(t_{0},x_{0})\in\mathbb{T}\times\mathbb{R}^{d}$,
$\sup_{k\in\mathbb{N}}\|v^{n_{k}}\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(D_{1}(t_{0},x_{0}))}\leq A_{0},\quad\sup_{(t,x)\in[t_{0},\infty)\times\mathbb{R}^{d},a\in a}|r(t,x,a)|+\sup_{t\in[t_{0},\infty)}|\delta(t)|A^{*},$
where $A_{0}$ is a finite constant independent of $A^{*}$. It follows from (2.1) and (3.2) that
$\lim_{t\to\infty}\sup_{k\in\mathbb{N}}\|v^{n_{k}}\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}([t,\infty)\times\mathbb{R}^{d})}=0.$
Moreover, by part (i), we can also obtain that,
$\|v^{\infty}\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(D_{1}(t_{0},x_{0}))}\leq A_{0},\quad\sup_{(t,x)\in[t_{0},\infty)\times\mathbb{R}^{d},a\in a}|r(t,x,a)|+\sup_{t\in[t_{0},\infty)}|\delta(t)|A^{*}.$
In conclusion,
$\lim_{t\to\infty}\|v^{\infty}\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}([t,\infty)\times\mathbb{R}^{d})}=0.$ (4.8)
Step 2. We then show $V^{\pi^{\infty}}(t,x)=v^{\infty}(t,x)$ for any $(t,x)\in\mathbb{T}\times\mathbb{R}^{d}$. Fix an arbitrary $(t,x)\in\mathbb{T}\times{\mathbb{R}^{d}}$. By the fact $b^{\pi^{\infty}}\in L^{\infty}({\mathbb{R}^{d}})$ and conditions on $\sigma$ stated in Assumption 2.2, the SDE
@@
admits a unique strong solution (see [26, Theorem 1]) with a density function belonging to $L^{q}([0,T]\times\mathbb{R}^{d})$ for any $T>0$ with $q=\frac{d+2}{d+1+\alpha}=\frac{p}{1-p}$ (see, e.g., [23, Theorem 9.1.9]). [^1] Here the result provided in [23, Theorem 9.1.9] is more general that the density belongs to $L^{q}(\mathbb{T}\times\mathbb{R}^{d})$ for all $q>1$. Given an arbitrary $\varepsilon>0$, thanks to (2.1) and (4.8), there exits some $T_{0}>0$ such that
$\left|E_{x}\left[\int_{T-t}^{\infty}r^{\pi^{\infty}}(t+s,X^{\pi^{\infty}}_{s})d\right]\right|+\|v^{\infty}\|_{L^{\infty}([T,\infty)\times\mathbb{R}^{d})}<\varepsilon/3\quad\forall T\geq T_{0}.$ (4.9)
@@
$\leq\|v^{\infty}\|_{L^{\infty}(\mathbb{T}\times\mathbb{R}^{d})}\mathbb{P}_{x}(\rho_{N}\leq T_{0})+\|v^{\infty}\|_{L^{\infty}([T,\infty)\times\mathbb{R}^{d})}\leq A^{*}\frac{\varepsilon}{3A^{*}}+\frac{\varepsilon}{3}=\frac{2\varepsilon}{3},\quad\forall N\geq N_{0}\vee T_{0}.$
Take any $N\in\mathbb{N}$ with $N\geq T\vee N_{0}$. In view of $v^{\infty}\in W^{1,2}_{p}(D_{N})\cap{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(D_{N})$, we can apply Itô-Krylov formula (see e.g. [16, Section 2.10, Theorem 1]) to $v^{\infty}(t+s,X^{\pi^{\infty}}_{s})$ to get
@@
Denote $p^{x}(s,y):=\mathbb{P}(X^{\pi^{\infty}}_{s}\in dy|s\leq\rho_{N})$. Taking expectations on both sides leads to
@@
$\qquad\qquad+\frac{1}{2}\operatorname{tr}\left((\sigma\sigma^{T})(X^{\pi^{\infty}}_{s})D^{2}_{x}v^{\infty}(t+s,X^{\pi^{\infty}}_{s})\right)\bigg)ds$

### 04 Apr 2026 v2

@@ 3 Existence of Equilibrium under Entropy Regularization
$\|\bar{V}^{n}\|_{{\mathcal{C}}^{1,2}_{\alpha/2,\alpha}(D_{N_{0}})}\leq C\left(\|\bar{V}^{n}\|_{L^{\infty}(D_{N_{0}+1})}+\|f^{n}\|_{{\mathcal{C}}_{\alpha/2,\alpha}(D_{N_{0}+1})}\right),$ (3.49)
@@
We are now ready to prove Theorem 3.1 Part (iii).
Proof of Theorem 3.1 Part (iii):.
Fix an entropy weight $0<\lambda\leq\lambda_{0}$ and an arbitrary Hölder constant $0<\beta<\alpha$. Lemmas 3.3 and 3.5 show that the mapping $\Phi_{\lambda}$ defined in (3.32) is a continuous mapping from the compact convex set ${\mathcal{M}}_{\lambda}$, equipped with norm $\|\cdot\|_{{\mathcal{C}}^{1,2,\operatorname{\textbf{wg}}}_{\beta/2,\beta}(\mathbb{T}\times\mathbb{R}^{d})}$, to itself. Hence, by Schauder fixed-point theorem (see, e.g., [22, Theorem 5.28]), $\Phi_{\lambda}$ has a fixed point $w\in{\mathcal{M}}_{\lambda}$ such that $w=V^{\pi^{*}}_{\lambda}$ with $\pi^{*}(x,a)=\Gamma_{\lambda}(x,D_{x}w(0,x),a)$. As a result, Lemma 3.3 yields that $V^{\pi^{*}}_{\lambda}$ satisfies (3.6). By taking $t=0$ in (3.19), together with the definition of $\Gamma_{\lambda}$ in (3.12), we see that $V^{\pi^{*}}_{\lambda}$ also satisfies (3.5). Then, we can conclude from Theorem 3.1 Part (ii) that $\pi^{*}$ is a regularized equilibrium with the entropy parameter $\lambda$. Moreover, the estimates of $V^{\pi^{*}}_{\lambda}$ are direct consequences from $V^{\pi^{*}}_{\lambda}\in{\mathcal{M}}_{\lambda}$ and Lemma 3.3. ∎
Remark 3.2.
We emphasis that, the existence of a classical solution to the EEHJB system (3.5)–(3.6) (i.e., the existence of an equilibrium under entropy regularization) stated in Theorem 3.1 Part (iii) is not restricted to the case of $\lambda\in(0,\lambda_{0}]$, but holds valid for all $\lambda>0$ by simply allowing $A^{*}$ to depend on $\lambda$. First, the sub-linear growth estimate in Lemma 3.1 holds for any $\lambda>0$. Indeed, as stated in [1, Lemma 1], for an arbitrary $\lambda>0$,
$|\mathcal{H}(\Gamma_{\lambda}(x,p,a))|=\left|\int_{U}\ln(\Gamma_{\lambda}(x,p,a))\Gamma_{\lambda}(x,p,a)da\right|\leq\kappa+\ell\ln(1+|p|),\quad\forall y\in\mathbb{R}^{d},$
where $\kappa>0$ depends on $\lambda$, $\ell$, $\operatorname{\texttt{Leb}}(U)$ and the parameters in Assumption 3.1. When $\lambda\leq\lambda_{0}$, $\kappa$ can be further refined as $\kappa=K_{6}+K_{7}|\ln\lambda|$ with $K_{6},K_{7}$ independent of $\lambda$. Thus, for any $\lambda>0$, the fixed point argument in this section remains valid.
In the analysis above, we have deliberately chosen the estimates of the value function $V^{\pi}$ to be independent of $\lambda$ because these estimates serve as a key tool in the convergence analysis of the next section as $\lambda{\to}0$.
@@ 4 Existence of Equilibrium by Vanishing Entropy Regularization
In this section, we return to the original time-inconsistent control problem without entropy regularization. To address the existence of equilibrium, we aim to establish a core convergence result: as the entropy parameter $\lambda{\to}0$, solutions of the EEHJB system converge to a solution to the original EHJB system in a proper sense such that the limit of the regularized equilibrium is a relaxed equilibrium in the original problem.
The next main result gives a sufficient condition, new to the literature, for the existence of an equilibrium based on the existence of a strong (rather than classical) solution to the EHJB system.
Theorem 4.1 (A sufficient condition for equilibrium).
For a measurable strategy $\pi^{*}:\mathbb{R}^{d}\to{\mathcal{P}}(U)$, if $V^{\pi^{*}}\in{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})\cap W^{1,2,\operatorname{\textbf{ul}}}_{p}(\mathbb{T}\times\mathbb{R}^{d})$ with $\sup_{x\in\mathbb{R}^{d}}\|V^{\pi^{*}}(\cdot,x)\|_{{\mathcal{C}}^{1}_{\alpha/2}(\mathbb{T}\times\mathbb{R}^{d})}<\infty$, and $V^{\pi^{*}}$ is a strong solution to the EHJB system
$+\sup_{\varpi\in{\mathcal{P}}(U)}\left\{\int_{U}\left[b(x,a)D_{x}u(0,x)+r(0,x,a)\right]\varpi(da)\right\}\;\text{a.e on $\mathbb{R}^{d}$},$ (4.1)
$0=$ $\partial_{t}u(t,x)+\frac{1}{2}\operatorname{tr}\left((\sigma\sigma^{T})(x)D^{2}_{x}u(t,x)\right)+b^{\pi^{*}}(x)D_{x}u(t,x)+r^{\pi^{*}}(t,x)\;\text{a.e. on $\mathbb{T}\times\mathbb{R}^{d}$},$ (4.2)
then $\pi^{*}$ is an equilibrium in Definition 2.1.
Next, under our previous model assumptions, we further show that the existence of a strong solution to the EHJB system (4.1)–(4.2) is always guaranteed, thereby ensuring the existence of equilibrium in the original time-inconsistent control problem even when the classical solution of the EHJB system is unavailable.
Theorem 4.2 (Existence of a strong solution to EHJB system).
There exists a measurable strategy $\pi^{*}:\mathbb{R}^{d}\to{\mathcal{P}}(U)$ such that $V^{\pi^{*}}\in{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})\cap W^{1,2,\operatorname{\textbf{ul}}}_{p}(\mathbb{T}\times\mathbb{R}^{d})$ with $\sup_{x\in\mathbb{R}^{d}}\|V^{\pi^{*}}(\cdot,x)\|_{{\mathcal{C}}^{1}_{\alpha/2}(\mathbb{T}\times\mathbb{R}^{d})}<\infty$, and $V^{\pi^{*}}$ is a strong solution to the EHJB system (4.1)–(4.2), and hence $\pi^{*}$ is an equilibrium in Definition 2.1 by Theorem 4.1.
In the following subsections, we first prove Theorem 4.2 by showing that the classical solutions to the EEHJB system will converge to a strong solution to the EHJB system when the entropy regularization vanishes. Then, by leveraging the convergence results, we will further develop some new verification arguments to etablish the weaker sufficient condition for the existence of equilibrium in Theorem 4.1.
4.1 Proof of Theorems 4.2
Let us take a sequence $(\lambda_{n})_{n\in\mathbb{N}}$ with $\lambda_{n}\to 0+$. For each $n\in\mathbb{N}$, Theorem 3.1 Part (iii) states the existence of a regularized equilibrium $\pi^{n}$ with entropy parameter $\lambda_{n}$ in Gibbs form such that $\|V^{\pi^{n}}_{\lambda_{n}}\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(\mathbb{T}\times\mathbb{R}^{d})}\leq A^{*}$ for all $n\in\mathbb{N}$. Let us further denote $v^{n}(t,x):=V^{\pi^{n}}_{\lambda_{n}}(t,x)$ for each $n$. It then follows from Theorem 3.1 Part (i) that $(v^{n},\pi^{n})$ satisfies
@@
And we will construct a strong (instead of classic) solution to the EHJB system (4.1)–(4.2) by searching a subsequence limit of $(v^{n},\pi^{n})_{n\in\infty}$.
@@
There exist a subsequence $(v^{n_{k}},\pi^{n_{k}})_{k\in\mathbb{N}}$ and a pair $(v^{\infty},\pi^{\infty})$ satisfying the following.
@@
    $\lim_{k\to\infty}\|v^{n_{k}}-v^{\infty}\|_{{\mathcal{C}}^{0,1}_{\beta/2,\beta}(D_{N})}=0\quad\forall 0<\beta<\alpha,$
@@
    $\pi^{\infty}:\mathbb{R}^{d}\to{\mathcal{P}}(U)$ is Borel measurable, and $\pi^{n_{k}}$ converges to $\pi^{\infty}$ in the sense that
@@
    for any test function $\phi(x,a):\mathbb{R}^{d}\times U\to\mathbb{R}^{d}$ that is continuous on the control variable $a$ for each $x$ and satisfies $\int_{\mathbb{R}^{d}}\max_{a}|\phi(x,a)|dx<\infty.$ Moreover,
[…]
@@
    $V^{\pi^{\infty}}=v^{\infty}$, $\sup_{x\in\mathbb{R}^{d}}\|v^{\infty}(\cdot,x)\|_{{\mathcal{C}}^{1}_{\alpha/2}(\mathbb{T})}\leq A^{*}$.
@@
where $A^{*}$ is the finite constant stated in Theorem 3.1, and $A_{N}=C(N^{d+1})^{1/p}$ with $C$ being a finite constant independent of $d,p$ and $N$. For each $N$, define
@@
Since $\sup_{n\in\mathbb{N}}\|v^{N-1,n_{k}}\|_{W^{1,2}_{p}(D_{N})}\leq A_{N}A^{*}$, by Banach–Alaoglu theorem, we can extract a subsequence of $(v^{N-1,n_{k}})_{k\in\mathbb{N}}$, denoted by $(v^{N,n})_{n\in\mathbb{N}}$, such that, for any test function $\phi\in L_{q}(D_{N})$,
$\lim_{n\to\infty}\int_{D_{N}}\phi(t,x)\partial_{t}v^{N,n}(t,x)dtdx=\int_{D_{N}}\phi(t,x)w(t,x)dtdx,$
$\lim_{n\to\infty}\int_{D_{N}}\phi(t,x)\partial^{2}_{x_{i}x_{j}}v^{N,n}(t,x)dtdx=\int_{D_{N}}\phi(t,x)w^{ij}(t,x)dtdx,\quad\forall 1\leq i,j\leq d,$
where the functions $w,w^{ij}$ belong to $L_{p}(D_{N})$. Then for any test function $\phi\in C^{\infty}_{c}(D_{N})$, $\partial_{t}\phi,\partial_{x_{i}}\phi\in L_{q}(D_{N})$ and for each $1\leq i,j\leq d$,
$-\int_{D_{N}}\phi(t,x)w^{ij}(t,x)dtdx=$ $-\lim_{n\to\infty}\int_{D_{N}}\phi(t,x)\partial^{2}_{x_{i}x_{j}}v^{N,n}(t,x)dtdx$
$=$ $\lim_{n\to\infty}\int_{D_{N}}\partial_{x_{i}}\phi(t,x)\partial_{x_{j}}v^{N,n}(t,x)dtdx=\int_{D_{N}}\partial_{x_{i}}\phi(t,x)\partial_{x_{j}}u^{N}(t,x)dtdx.$
Thus, $w^{ij}$ is the weak derivative $\partial_{x_{i}}(\partial_{x_{j}}u^{N})$ for each $1\leq i,j\leq d$. A similar argument shows that $w$ is the weak derivative $\partial_{t}u^{N}$. Therefore, $u^{N}\in W^{1,2}_{p}(D_{N})$, and
@@
As a result, (4.7) and the following holds for the subsequence $(v^{N,n})_{n\in\mathbb{N}}$ and $u^{N}$
@@
By taking the diagonal subsequence $(v^{N,N})_{N\in\mathbb{N}}$, we obtain for each $K\in\mathbb{N}$, any $l\in\mathbb{N}$ and any multi-index $a$ satisfying $0\leq 2l+|a|_{l^{1}}\leq 2$,
$\lim_{N\to\infty}\|v^{N,N}-v^{\infty}\|_{{\mathcal{C}}^{0,1}_{\beta/2,\beta}(D_{K})}=0,\quad\forall 0<\beta<\alpha,$
@@
Since for a generic function $g\in L^{p}(D_{1}(t_{0},x_{0}))$,
$\|g\|_{L^{p}(D_{1}(t_{0},x_{0}))}=\sup_{\|\phi\|_{L^{q}(D_{1}(t_{0},x_{0}))}\leq 1}\int_{D_{1}(t_{0},x_{0})}\phi(t,x)g(t,x)dtdx,$
for any test function $\phi$ with $\|\phi\|_{L^{q}(D_{1}(t_{0},x_{0}))}\leq 1$, we have
@@
$\leq$ $\liminf_{N\to\infty}\|\partial_{t}D^{a}_{x}v^{N,N}\|_{L^{p}(D_{1}(t_{0},x_{0}))}.$
It follows that
$\|\partial_{t}D^{a}_{x}v^{\infty}\|_{L^{p}(D_{1}(t_{0},x_{0}))}\leq\liminf_{N\to\infty}\|\partial_{t}D^{a}_{x}v^{N,N}\|_{L^{p}(D_{1}(t_{0},x_{0}))}\leq A^{*}\quad\forall 0\leq 2l+|a|_{l^{1}}\leq 2,$
and hence, $\|v^{\infty}\|_{W^{1,2,\operatorname{\textbf{ul}}}_{p}(\mathbb{T}\times\mathbb{R}^{d})}\leq A^{*}$.
Part (ii). Next, we consider the associated sequence $(\pi^{N,N})_{N\in\mathbb{N}}$ with respect to the sequence $(v^{N,N})_{N\in\mathbb{N}}$. By Young measure theory (see, e.g., [29, Chapter IV]), a subsequence $(\pi^{n_{k}})_{k\in\mathbb{N}}$ of $(\pi^{N,N})_{N\in\mathbb{N}}$ and a Borel function $\pi^{\infty}:\mathbb{R}^{d}\to{\mathcal{P}}(U)$ exist such that the convergences stated in part (ii) holds. In conclusion, the corresponding subsequence $(v^{n_{k}},\pi^{n_{k}})_{k\in\mathbb{N}}$ and the pair $(v^{\infty},\pi^{\infty})$ indeed satisfy (4.6) and Part (i).
Moreover, for any test function $\phi\in L^{1}(\mathbb{R}^{d})$, by Assumption 3.1, $a\mapsto b(x,a)\phi(x)$ is Lipschitz continuous for all $x$ and $\int_{\mathbb{R}^{d}}\max_{a\in U}|b(x,a)\phi(x)|dx\leq\|b(x,a)\|_{L^{\infty}(\mathbb{R}^{d}\times U)}\int_{\mathbb{R}^{d}}\phi(x)dx<\infty$. As a result,
$\lim_{k\to\infty}\int_{\mathbb{R}^{d}}b^{\pi^{n_{k}}}(x)\phi(x)dx=$ $\lim_{k\to\infty}\int_{\mathbb{R}^{d}}\left(\int_{U}[b(x,a)\phi(x)]\pi^{n_{k}}(x,a)da\right)dx$
$=$ $\int_{\mathbb{R}^{d}}\left(\int_{U}[b(x,a)\phi(x)]\pi^{\infty}(x,da)\right)dx=\int_{\mathbb{R}^{d}}b^{\pi^{\infty}}(x)\phi(x)dx\quad\forall\phi\in L^{1}(\mathbb{R}^{d}).$
A similar argument also holds for $r(t,\cdot)$. As a result, (4.6) holds.
Part (iii). Now we show $V^{\pi^{\infty}}(t,x)=v^{\infty}(t,x)$ for any $(t,x)\in\mathbb{T}\times\mathbb{R}^{d}$. By Theorem 3.1 Part (iii),
$\sup_{k\in\mathbb{N}}\|v^{n_{k}}\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}([t,\infty)\times\mathbb{R}^{d})}\leq A^{*}\Psi(t).$
Then it follows from Part (i) that,
$\lim_{t\to\infty}\|v^{\infty}\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}([t,\infty)\times\mathbb{R}^{d})}\leq\lim_{t\to\infty}\sup_{k\in\mathbb{N}}\|v^{n_{k}}\|_{{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}([t,\infty)\times\mathbb{R}^{d})}=0.$ (4.8)
Fix an arbitrary $(t,x)\in\mathbb{T}\times{\mathbb{R}^{d}}$. By the fact $b^{\pi^{\infty}}\in L^{\infty}({\mathbb{R}^{d}})$ and conditions on $\sigma$ stated in Assumption 2.2, the SDE
@@
admits a unique strong solution (see [26, Theorem 1]) with a density function belonging to $L^{q}([0,T]\times\mathbb{R}^{d})$ for any $T>0$ with $q=\frac{d+2}{d+1+\alpha}=\frac{p}{1-p}$ (see, e.g., [23, Theorem 9.1.9]). [^1] The result in [23, Theorem 9.1.9] is more general that the density belongs to $L^{q}(\mathbb{T}\times\mathbb{R}^{d})$ for all $q>1$. Given an arbitrary $\varepsilon>0$, thanks to (2.1) and (4.8), there exits some $T_{0}>0$ such that
$\left|E_{x}\left[\int_{T-t}^{\infty}r^{\pi^{\infty}}(t+s,X^{\pi^{\infty}}_{s})d\right]\right|+\|v^{\infty}\|_{L^{\infty}([T,\infty)\times\mathbb{R}^{d})}<\frac{\varepsilon}{3}\quad\forall T\geq T_{0}.$ (4.9)
@@
$\leq\|v^{\infty}\|_{L^{\infty}(\mathbb{T}\times\mathbb{R}^{d})}\mathbb{P}_{x}(\rho_{N}\leq T_{0})+\|v^{\infty}\|_{L^{\infty}([T,\infty)\times\mathbb{R}^{d})}\leq A^{*}\frac{\varepsilon}{3A^{*}}+\frac{\varepsilon}{3}=\frac{2\varepsilon}{3}\quad\forall N\geq N_{0}\vee T_{0}.$
Take any $N\in\mathbb{N}$ with $N\geq T\vee N_{0}$. In view of $v^{\infty}\in W^{1,2}_{p}(D_{N})\cap{\mathcal{C}}^{0,1}_{\alpha/2,\alpha}(D_{N})$, we can apply Itô-Krylov formula (see e.g. [16, Section 2.10, Theorem 1]) to $v^{\infty}(t+s,X^{\pi^{\infty}}_{s})$ to obtain
@@
Denote $p^{x}(s,y):=\mathbb{P}_{x}(X^{\pi^{\infty}}_{s}\in dy|s\leq\rho_{N})$ and note that $p^{x}\in L^{q}(D_{N})$. [^2] Since $C:=\mathbb{P}^{x}(N\leq\rho_{N})$ is a finite value for a fixed $N$, and $\mathbb{P}^{x}(s\leq\rho_{N})\geq\mathbb{P}^{x}(N\leq\rho_{N})$ for all $s\in[0,s]$, we have $\int_{D_{N}}|p^{x}(d,y)|^{q}dsdy=\int_{D_{N}}\left|\frac{\mathbb{P}^{x}(X^{\pi^{\infty}}_{s}\in dy,s\leq\rho_{N})}{\mathbb{P}^{x}(s\leq\rho^{N})}\right|^{q}dsdy\leq\frac{1}{C^{q}}\int_{D_{N}}|\mathbb{P}^{x}(X^{\pi^{\infty}}_{s}\in dy)|^{q}dsdy<\infty$. Integrating both sides of above equality over $0,N\wedge\rho_{N}-t]$ and then taking expectations leads to
@@
$\qquad\qquad\qquad\qquad\qquad+\frac{1}{2}\operatorname{tr}\left((\sigma\sigma^{T})(X^{\pi^{\infty}}_{s})D^{2}_{x}v^{\infty}(t+s,X^{\pi^{\infty}}_{s})\right)\bigg)ds$